Find a so that the graph of f(x) = loga(x) contains the point (1/3, -6). Give your answer correct to 2 decimal places.
Answer by swincher4391(1107)

Basically, this question is asking a^-6 = 1/3 for what value of a satisfies this equation?
Well, to solve this we must *undo* a^-6 by raising it to the -1/6 power.
(a^-6)^(-1/6) = (1/3)^(-1/6) = 1.20 <---- Answer

To check, we put in log(1/3) / log(1.2) [change of base formula] and we get roughly -6, since we rounded our base to 1.20.
Thus a = 1.20
Hopefully this helped!
If this was not clear, please e-mail me at onlyyourmathtutor@yahoo.com.